OpenStudy (gabylovesyou):

Which of the following represents the linear equation 3x = 12 − 2y in standard form? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need to get the equation into Ax+By = C form

OpenStudy (anonymous):

Your standard form is : ax + by = c So, you would add 2y to both sides. That will produce: 3x + 2y = 12 − 2y + 2y 3x + 2y = 12

OpenStudy (gabylovesyou):

got u thanks :)

OpenStudy (anonymous):

uw!